OpenStudy (anonymous):

write a quadratic inequality for which the solution is -2

OpenStudy (anonymous):

f (x) = ax^2 + bx + c < 0 (x + 2)(x -5) = x^2 - 5x + 2x - 10 = x^2 - 3x - 10 < 0. The parabola is upward. Between the 2 real roots (-2) and (5), f(x) is negative as opposite in sign to a = 1. In this interval (-2, 5), the parabola is below the x-axis.

OpenStudy (anonymous):

so wait the answer is x^2-3x-10<0?

OpenStudy (jdoe0001):

to do that for say an equation, you'd use the values for "x" to get the roots, that is \(\bf x=-2\implies x+2=0\implies (x+2)=0 \\ \quad \\ x=5\implies x-5=0\implies (x-5)=0 \\ \quad \\ (x+2)(x-5)=0\)

OpenStudy (jdoe0001):

in this case is an inequality, so you could write it as \(\bf (x+2)(x-5)<0\qquad (x+2)(x-5)>0\)

OpenStudy (anonymous):

which is euqal to x^2-3x-10>0 right?

OpenStudy (jdoe0001):

either one so the quadratic equation will be the multiplication of both of the roots yes, which is equals to \(\bf (x+2)(x-5)<0\implies x^2-3x-10<0 \\ \quad \\ (x+2)(x-5)>0\implies x^2-3x-10>0\) so use the form that gives you the values back as you need them

OpenStudy (jdoe0001):

for this case, it'd be the 1st one, because \(\bf (x+2)(x-5)<0\implies x^2-3x-10<0 \\ \quad \\ x^2-3x-10<0\implies (x+2)(x-5)<0\to \begin{cases} x+2<0\implies x<-2 \\ \quad \\ \bf x-5<0\implies x<5 \end{cases}\)

OpenStudy (anonymous):

ok i think i get it :D thanks, do u know alot about like radians and stuff?

OpenStudy (jdoe0001):

just post anew, thus more exposure, and we can revise each other

OpenStudy (anonymous):

ok thanks :D

OpenStudy (jdoe0001):

hmm just a quick revision \(\bf x+2>0\implies x>-2 \\ \quad \\ x<5\implies x-5<0\implies -x-5>0 \\ \quad \\ (x+2)(-x-5)>0\)

OpenStudy (jdoe0001):

rather \(\bf x+2>0\implies x>-2 \\ \quad \\ x<5\implies x-5<0\implies -x+5>0 \\ \quad \\ (x+2)(-x+5)>0\)